OpenStudy (anonymous):

Peter measured the width of one cardboard strip as 2 3/4 ft. A second cardboard strip measured 3 2/3 ft in width. Estimate the combined wifth of the cardboard??

OpenStudy (bahrom7893):

11/4+11/3=(33+44)/12=77/12

OpenStudy (anonymous):

?